OpenStudy (anonymous):

Solve this system of equations using any method. 3x – y = 4 x + 2y = 13 A. (–1, 7) B. (4, 8) C. (5, 11) D. (3, 5) ~will give medal!~

OpenStudy (anonymous):

ok, give me a minute to work it out.

OpenStudy (anonymous):

it has that question and ur other ones. and answers/

OpenStudy (anonymous):

I cant figure it out, just listen to siega. :) sorry.

OpenStudy (anonymous):

thats not the right questions though there wrong i went to go look at it. And there all the wrong questions

OpenStudy (anonymous):

\[3x – y = 4\]\[ x + 2y = 13 \] multiply the first equation by 2, get \[6x-2y=8\]\[x+2y=13\] add to get \[7x=21\] so \(x=3\)

OpenStudy (anonymous):

So is it d?

OpenStudy (anonymous):

i don't know i didn't look

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

D. (3, 5)

OpenStudy (anonymous):

yeah, it is D